SIPB Office Spare IP Addresses

/afs/sipb/admin/office/office-laptops.tex November 29, 2006

These addresses are extra office IP addresses intended to be used for things like new virtual machines, for machines that are rarely on like the alpha and sgi in the office rack, for giving laptops static IPs in the office, etc. Please mark this document if you give one of these IPs a specific purpose (like a virtual machine running a service).

• mega-man (18.187.2.216)

• kid-icarus (18.187.2.218)

• general-dink (18.187.1.79)

• planet-zorp (18.187.1.84)

• kitzi (18.187.1.152) [sipb-svn vm] Gateway: 18.187.0.1 Netmask: 255.255.0.0 Nameserver: 18.70.0.160

18.71.0.151 18.72.0.3